OpenStudy (4meisu):

Each day a runner trains for a 10km race. On the day she runs 1000 m, and then increases the distance by 250m on each subsequent day. On which day does she run a distance of 10km in training? What is the total distance she will have to run in training by the end of that day?

OpenStudy (jamesj):

Let s(x) be the distance she runs on day x. Then s(1) = 1000 s(2) = 1250 s(3) = 1500 ... What is the general formula for s(x)?

OpenStudy (jamesj):

s(x) = 1000 + .... what?

OpenStudy (jamesj):

still there?

OpenStudy (4meisu):

Yeah sorry, the one I can find from the textbook is s(x) a1(1-r^n)/1-r

OpenStudy (jamesj):

That formula is completely wrong here. Try and follow the logical path I am leading you on.

OpenStudy (jamesj):

s(1) = 1000 + 0 s(2) = 1000 + 250 s(3) = 1000 + 500 s(4) = 1000 + 750 ... See what the formula is now for s(x)? s(x) = 1000 + (something involving x)

OpenStudy (jamesj):

I'll tell you then. s(x) = 1000 + 250(x-1) This is the formula for an arithmetic sequence. Make sense?

OpenStudy (4meisu):

So far, yes.

OpenStudy (jamesj):

If so, the way to answer the first question, "On which day does she run a distance of 10km in training?" is to ask for what value of x is s(x) = 10,000 m ?

OpenStudy (jamesj):

So solve now for x.

OpenStudy (4meisu):

Okay, thanks!